OpenStudy (anonymous):

Find a quadratic function given by C(x) = a(x - 1959)2 + k that models the data

OpenStudy (anonymous):

(1959,300), (1982,2000), (1990,4000)

OpenStudy (asnaseer):

I assume your expression is:\[C(x)=a(x-1959)^2+k\]

OpenStudy (asnaseer):

substitute the first pair of values into this, i.e. you are given C(x)=300 when x=1959

OpenStudy (asnaseer):

that will allow you to calculate the value of k. then substitute the next pair of values in - i.e. C(x)=2000 when x=1982 and that should allow you to calculate the value for 'a'.

OpenStudy (anonymous):

so c(300)=a(1959-1959)^2+k =a(0)+k k=300

OpenStudy (asnaseer):

you have the value for k correct, but should have written as follows: C(1959) = 300 = a1959-1959)^2 + k

OpenStudy (asnaseer):

missed a bracket: C(1959) = 300 = a(1959-1959)^2 + k

OpenStudy (anonymous):

c(1982)=2000=a(1982-1959)^2+k =a(23)^2+k i am stuck here

OpenStudy (asnaseer):

you know the value for k, so put that in

OpenStudy (anonymous):

oh, gotcha. So, c(1982)=2000=a(1982-1959)^2+k =a(23)^2+300 2000=a(23)^2+300 =529a+300-2000 a=-1700/529 so a=3.21?

OpenStudy (asnaseer):

almost, your steps have some mistakes in them, it should be: 2000 = a(23)^2 + 300 2000 - 300 = 529a 1700 = 529a a = 1700 / 529 = 3.21

OpenStudy (anonymous):

so a will change for every c(x)?

OpenStudy (asnaseer):

no - if you use the values of 'a' and 'k' that you found an re-write the expresion as:\[C(x)=3.21(x-1959)^2+300\]then you should find that this gives you the correct value for C(x) for every x value given to you in the question. in other words, the values you found for 'a' and 'k' have given you a quadratic equation that satisfies all the data points.